The number of all possible triplets (a1, a2,a3) such that a1 + a2 cos2x +a3 sin 2 x = 0 for all x is


(a) 0     (b) 1   (c)3    (d) infinity


 

Grade:12

1 Answers

Ramesh V
14 years ago

A + B(1-t2) + Ct2      where t2 = sin2x   , i.e., 0 =< t2 =< 1 as -1<sin x<1

t2 = (2B-C)/(A+B)

0 =< (2B-C)/(A+B) =< 1

i.e., 2B >= C  and A+C>=B

the possible triplets for which the above inequalities satisfy are :

(A,B,C) in order of (0,1,1) (1,1,0) (0,0,0) (2,3,1) and so on

answer is infinity

--

regards

Ramesh